OpenStudy (anonymous):

At an amusement park, each premium ride costs $12 and each gold star ride is $8. Stan spends $96 on tickets and he gets 3 more premium rides than gold star rides.

OpenStudy (anonymous):

pAt an amusement park, each premium ride costs $12 and each gold star ride is $8. Stan spends $96 on tickets and he gets 3 more premium rides than gold star rides. How much does he spend on gold star rides?

OpenStudy (anonymous):

Ari is taking a group of people to the movies. Each child costs $4 and each adult is $9. How much will he pay if the group consists of 3 adults and 4 children?

OpenStudy (anonymous):

1. p = premium rides g = gold star ride 12p + 8g = 96 p = 3 + g now sub 3 + g in for p in the 1st equation 12(3 + g) + 8g = 96 -- distribute the 12 through the parenthesis 36 + 12g + 8g = 96 -- combine like terms 36 + 20g = 96 -- subtract 36 from both sides 20g = 96 - 36 20g = 60 -- divide both sides by 20 g = 3 now sub 3 in for g in the 2nd equation p = 3 + g p = 3 + 3 p = 6 Lets check it... 12p + 8g = 96 12(6) + 8(3) =96 72 + 24 = 96 96 = 96 (correct) She rides 3 gold star rides at $8 per ride...3 x 8 = $24 <--answer 2. Each child costs $4 and each adult cost $9. 3 adults would cost (3 x 9) = $27 4 children would cost (4 x 4) = $16 Total cost would be (27 + 16) = $43 <--answer